# Exchangeability in Neural Network Architectures and its Application to Dynamic Pruning

### v1

@@ 3.3 Exchangeable Parameters and Values in Popular Neural Networks
Normalization Layers. Normalization layers (NL) such as batch normalization BatchNorm and layer normalization LayerNorm are popular for stabilizing training. It applies a channel-wise scaling and bias. In CNNs, one normalization layer is usually inserted right before or after each activation function. We present an analysis for the latter case but the analysis also applies to the former case. Denote the NL functions as $\mathrm{NL}$ with parameters $\mathrm{nl}$, and $\mathrm{NL}_{i}$ denotes the NL applied to the $i$-th channel, parameterized by $\mathrm{nl_{i}}$. We instantiate $n=C_{2}$, $\zeta{}_{i}=W^{\prime}_{i\cdot}\oplus{}\mathrm{nl}_{i}\oplus{}W_{\cdot i}$, and $\xi{}_{i}=\mathrm{conv}(\mathrm{NL}_{i}(\mathrm{\sigma}{}(\mathrm{conv}(A,W^{\prime}_{i\cdot}))),W_{\cdot i})$.
@@
Embeddings Embeddings are used for many tasks such as NLP word2vec ; glove and graphs GCNWelling . Embeddings are “distributed” representations as the relevant information is represented in many dimensions. Our theory characterizes the symmetry often present in the embedding dimensions as statistically exchangeable. For example, in Word2vec word2vec , the NN architecture is simply an embedding layer (look-up table) and a fully connected layer, followed by a softmax operation $\mathrm{\sigma}{}$. It takes a word as input, and output likelihoods for each word to be in its context. Let $m$ be the number of words, $n$ be the embedding dimension, $A$ be the embedding matrix for all the words, $M$ be the fully connected layer weight matrix (both of size $m\times n$). The NN takes into input a word index $k$, and outputs $\mathrm{\sigma}{}(MA_{k}^{T})$. We instantiate $\zeta{}_{i}=A_{\cdot i}\oplus{}M_{\cdot i}$ and $\xi{}_{i}=A_{ki}M_{\cdot i}$. Alternatively, $\xi{}_{i}=A_{ki}$, which indicates that learned embedding dimensions are exchangeable.
Transformers Transformers are widely used in LLMs attention , vision visiontransformers , etc. We next analyze a decoder-only transformer architecture due to its popularity in recent LLMs llama ; gpt2 ; qwen . We formalize the simple case of single-head attention followed by a fully connected layer. This basic analysis can be extended to include normalization layers and skip connections using the techniques introduced above for CNNs.
The key insight here is that the output dimensions of the attention layer are exchangeable. Denote the input sequence length as $m$. Let the vector dimensions for each token be $d_{1},d_{2},d_{3}$ before attention, after attention, and after fully-connected layer respectively. The input $X$ has shape $d_{1}\times m$, the key, query, value matrices $K,Q,V$ all have shape $d_{2}\times d_{1}$. The fully connected layer’s weight $W$ has shape $d_{3}\times d_{2}$. The function of these two layers is $WVX\mathrm{\sigma}{}((QX)^{T}(KX))$, where $\mathrm{\sigma}{}$ is column-wise softmax. We instantiate $n=d_{2}$ and $\zeta{}_{i}=V_{i\cdot}\oplus{}W_{\cdot i}$ (or $\zeta{}_{i}=K_{i\cdot}\oplus{}Q_{i\cdot}\oplus{}V_{i\cdot}\oplus{}W_{\cdot i}$). Permuting $\zeta{}_{i}$’s is a parameter space symmetry because $(WP^{T})(PV)X\mathrm{\sigma}{}((QX)^{T}(KX))=WVX\mathrm{\sigma}{}((QX)^{T}(KX))$, for any permutation matrix $P$. The exchangeable values here are the attention layer’s output $\xi{}_{i}=V_{i\cdot}X\mathrm{\sigma}{}((QX)^{T}(KX))$, or the final output $\xi{}_{i}=W_{\cdot i}V_{i\cdot}X\mathrm{\sigma}{}((QX)^{T}(KX))$. Interestingly, this indicates that the intermediate activations in every decoder layer are exchangeable.
4 ExPrune Algorithm
We next present ExPrune algorithm, a dynamic pruning algorithm for NNs that leverages the exchangeability property identified in Section 3. Exchangeable values are identically distributed. Intuitively, they may contain overlapping information, which is a form of redundancy. In this section, we present our algorithm ExPrune that removes this redundancy on a per-input basis at inference, by terminating computation of exchangeable values early, when certain confidence metric is satisfied.
$\mathrm{confident}$, $h$, $n$
$\xi{}^{\prime}=<>$
for $i$ in $\mathbf{argsort}(h)$ do
     Compute $\xi{}_{i}$, $\xi{}^{\prime}=\xi{}^{\prime}\oplus{}\xi{}_{i}$
     if $\mathrm{confident}(\xi{}^{\prime})$ then
         return $\rho(\xi{})$
return $\rho(\xi{})$
(a) ExPrune general algorithm.
$\mathrm{negative\_predict}$, $\mathrm{h}$, $w$, $b$, $n$
$\xi{}^{\prime}=<>$, $n^{\prime}=0$
for $i$ in $\mathbf{argsort}(h)$ do
     Compute $\xi{}_{i}$, $\xi{}^{\prime}=\xi{}^{\prime}\oplus{}\xi{}_{i}$, $n^{\prime}$ += $1$
     if $\mathrm{negative\_predict}(\xi{}^{\prime},n^{\prime},w,b)$ then
         return 00
[…]
(b) ExPrune algorithm instantiated for neuron-level early negative prediction of ReLU.
Figure 3: Pseudocode for ExPrune algorithm. $\xi{}_{1},\ldots,\xi{}_{n}$ are exchangeable values to be computed. $h$ is heuristic prioritization scores, and by default $h_{i}=i$. $\rho$ is invariant to input permutations and can be approximately evaluated with fewer than $n$ input $\xi{}_{i}$’s.
4.1 General Dynamic Pruning Algorithm
Assume we want to compute a NN sub-module with function $\rho(\xi{})$, where $\xi{}_{i}$’s are exchangeable values. The input to the sub-module have been computed. $\rho$ is invariant to permutations of $\xi{}_{i}$’s, and can be approximated by $n^{\prime}$ ($n^{\prime}\leq n$) different $\xi{}_{i}$’s as input. Larger $n^{\prime}$ leads to more accurate approximation. Assume that we have a confidence metric $\mathrm{confident}$ that takes a set of $\xi{}_{i}$’s as input, and outputs whether we are confident about the current result. We can also optionally have a heuristic function $h$, where $h_{i}$ assigns a prioritization score for computation of $\xi{}_{i}$. For example, $h$ can be based on the cost to compute $\xi{}_{i}$, or the estimated importance of $\xi{}_{i}$.
Algorithm 3(a) presents the ExPrune general algorithm. It sequentially computes $\xi{}_{i}$’s in the order specified by $h$, and terminate early when the current result satisfies the confidence metric. Note that ExPrune can also optionally do confidence test less often to reduce overhead. This simple algorithm captures the idea behind many existing algorithms, e.g., finding the best arm in multi-armed bandit problem with smallest number of pulls possible medoids , adaptively computing the top k softmax values without full evaluation adaptivesoftmax , and channel-level dynamic pruning in CNNs RuntimeNeuralPruning .
4.2 Early Negative Prediction for ReLU
We present an specific instantiation of ExPrune that dynamically prunes computation for each ReLU-activated neuron. The permutation-invariant function $\rho$ is instantiated as summation. The algorithm exploits the property of ReLU that it outputs zero for negative input. We can therefore terminate computation of $\xi{}_{i}$’s when we predict based on the computed $\xi{}_{i}$’s that the final sum is likely negative. Note that this is only possible because $\xi{}_{i}$’s are exchangeable and thus identically distributed. In some cases, the sum of $\xi{}_{i}$’s is not directly processed by ReLU, but is first scaled and added biases, e.g., normalization layers, layer biases, shortcut connections. The negative prediction takes into account computed $\xi{}_{i}$’s, the number of $\xi{}_{i}$’s processed $n^{\prime}$, the scaling weight $w$, and the added bias $b$ when determining confidence. We devise two negative prediction methods as follows (assuming $w=1,b=0$ for simplicity).
  • Threshold. We set a predetermined threshold $T$ and terminate when the current mean is below threshold $\sum_{i=1}^{n^{\prime}}\xi{}^{\prime}_{i}<n^{\prime}T$. This simple method only requires one more operation per invocation, as the running sum is computed by NN already.
  • StatsTest. We perform a Wald’s test waldsequentialtest with confidence level $\alpha$, checking if
    $\Phi(\frac{\frac{1}{n}\sum_{i=1}^{n^{\prime}}\xi{}^{\prime}_{i}}{\sqrt{\frac{1}{n}\sum_{i=1}^{n^{\prime}}\xi{}^{\prime 2}_{i}-(\frac{1}{n}\sum_{i=1}^{n^{\prime}}\xi{}^{\prime}_{i})^{2}}})<\alpha,\mathrm{\ or\ equivalently\ }\frac{(\sum_{i=1}^{n^{\prime}}\xi{}^{\prime}_{i})^{2}}{n^{\prime}\sum_{i=1}^{n^{\prime}}\xi{}^{\prime 2}_{i}-(\sum_{i=1}^{n^{\prime}}\xi{}^{\prime}_{i})^{2}}<(\Phi^{-1}(\alpha))^{2}$
    where $\Phi$ is the cumulative density function of the standard normal distribution, and the right hand side of the simplified inequality can be stored as constant. This method introduces overhead that scales linearly the number of $\xi{}_{i}$’s, as it requires computing running sum of the squared term. It is more costly but potentially more accurate. Note that the assumptions of Wald’s test are not met by all exchangeable sequences, but we find that it works well in practice (Section 5). We discuss the assumptions in more detail in Section 6.
One parameter $T$ or $\alpha$ can be shared across many neurons, e.g., all neurons in a NN layer.
5 Evaluation
Task Dataset Models Fidelity Metric
Image Classification CIFAR10 cifar VGG11-BN vgg , ResNet18-BN ResNet Accuracy
Graph Property Prediction ogbg-molhiv MoleculeNet Graph Convolutional NN (GCN) GCNWelling ROC-AUC
Question Answering PIQA PIQA OPT-1.3B OPT (pretrained model) Accuracy
Table 1: Datasets and models. BN means the model is enhanced with BatchNorm BatchNorm .
5.1 Experimental Setup
We evaluate ExPrune for early negative prediction for ReLU (Section 4.2) on various models with ReLU activation functions. Table 1 summarizes our benchmarks, models, and metrics.
Datasets and Models. We have 3 datasets, covering 3 different tasks and input types. Please refer to the Appendix A for the details of dataset split. We use popular CNNs for image inputs, GCN for graph inputs, and pretrained transformer language model for text inputs. We choose OPT OPT because it is an off-the-shelf ReLU activated language model. Please refer to the Appendix A for the details of obtaining the trained model weights. We do early negative prediction for computation of exchangeable partial activations followed by ReLU, specifically, in all convolution layers in GCN, all convolution layers except for the first one in CNNs, and the first linear layer after each attention layer in OPT.
Baselines and Metrics. As noted in Section 2, there is no neuron-level dynamic pruning algorithm for modern NN architectures to the best of our knowledge. Therefore, we configure ExPrune with threshold and StatsTest, described in Section 4.2, and compare with the unoptimized inference. We configure ExPrune to perform one negative prediction when $n^{\prime}=32$ to reduce overhead, i.e., ExPrune invokes $\mathrm{negative\_predict}$ once when $n^{\prime}=32$ and terminate if confident, otherwise computes all other $\xi{}_{i}$’s. We choose $n^{\prime}=32$ as many statistics methods target sample sizes of at least 30 30samplesize . We use the default fidelity metrics for each dataset. We use the floating-point operation performed (FLOPs) as our performance metric, as it is a good proxy for inference efficiency relustrikeback . We take into account the overhead of ExPrune when calculating FLOPs. Threshold uses one FLOP as the threshold $n^{\prime}T$ is stored as a constant. StatsTest uses $62n^{\prime}+$ 6 FLOPs as it uses $2n^{\prime}$ FLOPs for computing the sum of $\xi{}_{i}^{\prime 2}$, and $6$ more to compute the Wald’s statistic and compare it to the stored threshold. We report FLOPs of the whole model for CNNs and GCNs as ExPrune is applied to most of the computation, and the FLOPs of all the linear layers in between an attention layer and a ReLU activation for OPT, as ExPrune is only applied to these layers. The FLOPs of these linear layers account for approximately $1/3$ of the total FLOPs in the unoptimized inference ding2024memoryformer ; OPT ; HuggingfaceTransformers .
Hyperparameter optimization. As different layers in NN may have different error sensitivities, we use Optuna optuna_2019 to find optimal parameter combinations for ExPrune. Specifically, each layer has one parameter $T$ (Threshold) or $\alpha$ (StatsTest). We tune the hyperparameters for 2000 trials on the validation set. Refer to Appendix A for Optuna configuration details. Using the data we have on validation set, we select a subset of promising parameter configurations to run on the test set. This emulates the process of selecting the hyperparameter combination for deployment. We iteratively select all the combinations on the fidelity-FLOPs Parato Frontier (dubbed one Parato slice), remove them from the set, and choose all on the next Parato slice. We include all points on the first 5 Parato slices.
5.2 Results and Analysis
(a) VGG11-BN
(b) ResNet18-BN
(c) GCN
(d) OPT (FLOPs number of linear layers.)
[…]
Performance of ExPrune. Figure 4 shows the performance of ExPrune. We use the default evaluation order ($h_{i}=i$). Across four models and compared to the unoptimized baseline, ExPrune is able to deliver 10.98–26.3% reduction in FLOPs with negligible ($<$0.1%) fidelity drop, and 21.01–39.05% reduction in FLOPs with at most 1% fidelity drop. We find that StatsTest performs much better than Threshold for CNNs. Though StatsTest has higher overhead, it offers more accurate negative prediction for CNNs. In GCN and OPT, StatsTest and Threshold perform similarly. This is because it takes fewer FLOPs to compute each exchangeable value $\xi{}_{i}$ in the 1D convolution of GCN and the linear layer of OPT, compared to 2D convolution in CNNs, and thus the overhead of StatsTest acounts for a larger portion in the total FLOPs.
(a) 78th Iteration (1.93% dense convolution layers)
(b) 79th Iteration (1.83% dense convolution layers)
(c) 80th Iteration (1.74% dense convolution layers)
(d) 78th, 79th, and 80th iterations
Figure 5: Fidelity-FLOPs scatter plots for statically pruned VGG11-BN models. FLOPs are normalized to largest model’s unoptimized baseline in (d). Colors and lines have the same meaning as in Figure 4.
ExPrune and Static Pruning. We demonstrate that ExPrune can be applied to statically pruned VGG11-BN models and offer additional reduction in FLOPs. We statically prune the VGG11-BN model by iteratively setting the 5% parameters in all convolution kernels with smallest magnitude to zero, and finetune the model on the training set to recover accuracy, following the practice of Song et al. SongPruning . Note that the static pruning also preserves the exchangeability similarly as training does (Section 3.2). We take the models at 78th, 79th, and 80th iterations, where only 1.93%, 1.83%, and 1.74% weights in convolution layers are left. We choose aggressively pruned models because we want to study how ExPrune works with already extremely compressed models. To work with these models, we configure ExPrune to let $h_{i}$ equals the number of FLOPs required to compute $\xi{}_{i}$ (number of none-zero weights in $\zeta{}_{i}$), computing the cheap $\xi{}_{i}$’s first. This corresponds to sorting channels of convolution kernels, which can be done statically and introduces no overhead during inference.
The results are shown in Figure 5. Across three pruned models, ExPrune still provides 10.24–11.11% reduction in FLOPs with negligible ($<$0.1%) accuracy drop, and 13.91–14.39% reduction in FLOPs with at most 1% accuracy drop, compared to the unoptimized inference on statically pruned models. Threshold achieves better results than StatsTest in pruned models, because in these models the exchangeable values $\xi{}_{i}$’s are cheaper to compute, making the overhead of StatsTest offset the FLOPs reduction of computing fewer $\xi{}_{i}$’s. Figure 5(d) shows all the points in three models compared together. We find that ExPrune combined with static pruning achieve better accuracy-performance trade-off than only static pruning. This indicates ExPrune composes with static pruning, because ExPrune is able to remove symmetry-induced redundancy, which cannot be removed by static pruning.
6 Limitations and Future Work
[…]
Other ExPrune Instantiations. We provided one ExPrune instantiation that targets ReLU activation, but ExPrune can work with other activation functions (e.g., GELU GELU and SiLU SiLU ) with adjusted confidence metric. It should be noted that ReLU can effectively replace GELU or SiLU for efficient inference relustrikeback . ExPrune can also be instantiated to other granularities (e.g., kernel, layer, branches) and target other forms of partial evaluation. For example, we describe another possible instantiation that works with a larger granularity in Appendix A.
ExPrune and Hardware Acceleration. ExPrune reduces FLOPs but also breaks certain structures of computation which are exploited in hardware accelerators. We note that SnaPEA snapea shared similar compute patterns to ExPrune, and it is demonstrated that the FLOP reduction of SnaPEA could translate to real speed up and energy saving with special architecture support. It is of great practical value to design special architectures and hardware support to accelerate ExPrune.
7 Conclusion
We present a novel theory that formalizes a type of symmetry in NN architectures as statistical exchangeability. We identify exchangeable parameters and intermediate values in popular NNs. Exploiting this insight, we devise a general dynamic pruning algorithm ExPrune that removes the redundancy induced by exchangeable intermediate values. We instantiate ExPrune for neuron-level dynamic pruning, by terminating the computation early when a negative input to ReLU is predicted. We demonstrate that ExPrune is able to provide large FLOPs reduction in image CNNs, GCNs, and LMs. In addition, we show that ExPrune is able to compose with static pruning, provding additional FLOPs reduction on CNNs that are heavily pruned statically.
Appendix A Technical Appendices and Supplementary Material
A.1 ExPrune Instantiated for Prediction Head in Classification
$\mathrm{winner\_confident}$, $h$, $n$
$\xi{}^{\prime}=<>$, $\mathrm{class\_scores}=<0,0,\ldots,0>$, $n^{\prime}=0$
for $i$ in $\mathbf{argsort}(h)$ do
     Compute $\xi{}_{i}$, $\xi{}^{\prime}=\xi{}^{\prime}\oplus{}\xi{}_{i}$, $\mathrm{class\_scores}$ += $\xi{}_{i}$, $n^{\prime}$ += 1
     $\mathrm{cur\_winner}$ = $\mathrm{argmax(class\_scores)}$
     if $\mathrm{winner\_confident}(\xi{}^{\prime},\mathrm{class\_scores},\mathrm{cur\_winner},n^{\prime})$ then
         return $\mathrm{cur\_winner}$
return $\mathrm{argmax(class\_scores)}$
Figure 6: Pseudocode for ExPrune algorithm instantiated for prediction head in classification. $\xi{}_{1},\ldots,\xi{}_{n}$ are exchangeable values to be computed. $h$ is heuristic prioritization scores, and by default $h_{i}=i$. $\xi{}_{i}$’s and $\mathrm{class\_scores}$ are in $\mathcal{R}^{k}$, where $k$ is the number of classes.
We present another possible instantiation of ExPrune for prediction head in NNs, i.e., the last fully-connected layer preceding the $\mathrm{argmax}$, which is widely used in modern NNs such as MLP, CNNs, GNNs, transformers for classification tasks. Note that this instantiation is not evaluated in this work as the prediction head often comprises only a small portion of the total computation. However, one can potentially adapt it to replace $\mathrm{argmax}$ with $\mathrm{softmax}$, and this sub-structure accounts for a large portion of computation in the attention computation in transformers. Let $k$ be the number of classes, each $\xi{}_{i}\in\mathcal{R}^{k}$ provides a partial score for each class. Note that the granularity of this instantiation is larger than the instantiation in Section 4.2, because each $\xi{}_{i}$ is a vector instead of a single value, i.e., it is the computation from one neuron to all $k$ output neurons, not to just $1$ output neuron. We provide two possible $\mathrm{winner\_confident}$ metrics as follows.
Threshold. We set a predetermined threshold and terminate when the margin of mean between the largest and second largest scores is below threshold, i.e., terminate when $c_{1}-c_{2}<nT$, where the largest and second largest values in $\mathrm{class\_scores}$ as $c_{1},c_{2}$, respective.
StatsTest. We can conduct a Wald’s test [70] for $\mathrm{cur\_winner}$ against each of other classes, and terminate when all $k-1$ tests pass. Each test compares the mean of the largest class score to another class score. Since multiple tests are involved, we can use the Holm-Bonferroni method [33] to assign adjusted confidence levels for each test, given the overall type-I error bound $\alpha$.
@@ A.2 Evaluation Details
Dataset split. The test set labels of PIQA [9] are not published, so we use the validation set as the test set, and $10\%$ of the training set as the validation set. This is reasonable because the pretrained OPT model was not trained on PIQA training set. For CIFAR10 [42], we use a fixed split of the training set, with 90% samples used for training, 10% used as validation set, and use the default test set. We use the default dataset splits for ogbg-molhiv dataset [73].
Model architectures. We use the default OPT architecture from HuggingFace [72]. The GCN architectures follow Kipf and Welling [39]. We use the adapted CNNs [62, 30] for CIFAR10 from https://github.com/kuangliu/pytorch-cifar. Specifically, the first layer and the pooling layer before the prediction head are adapted in size for the image size and class number in CIFAR10.
Obtaining trained models. We train image CNNs and GCNs locally using the training set, and used the pretrained weights for OPT [75] from HuggingFace [72]. For local training of CNNs, we use AdamW optimizer, $5\times 10^{-3}$ learning rate, $16$ batch size, $10^{-4}$ weight decay, 1cycle learning rate scheduler, and 30 epochs. When statically pruning VGG11-BN, we use the same training scheme in every pruning iteration to finetune the model after 5% parameters in all convolution kernels are set to zero. For local training of GCN, we use AdamW optimizer, $10^{-3}$ learning rate, $32$ batch size, $10^{-4}$ weight decay, 1cycle learning rate scheduler, and 80 epochs.
Optuna hyperparameter tuning. Across all models, we set the range of $\alpha$ as $[0,0.5]$ in StatsTest for all layers. The range of $T$ in Threshold is $[-30,0]$ for CNNs, $[-1,0]$ for GCN, $[-0.005,0]$ for OPT. For statically pruned models, we additionally add a parameter $r$ in range $[0.1,0.5]$ to tune for each layer, which controls when ExPrune is disabled. When the ratio of the total FLOPs that can be potentially pruned for a channel’s computation to the overhead of ExPrune is below $r$, ExPrune is disabled. For each model, we provide a set of initial points to warm up Optuna’s surrogate model. Please see our code for details.
A.3 Compute Resources
We run most of our evaluation on a local laptop (MacBook M3 Pro with 36GB memory with 12 cores). We run the hyperparameter tuning of the OPT model with a NVIDIA GeForce RTX 4090 GPU.
A.4 Broader Impact
ExPrune enables deploying AI with fewer resources, and makes AI more energy efficient and environment-friendly. It also facilitates AI inference on edge devices with limited resources, where data is collected and processed locally without sending it to centralized servers, enhancing user data security and privacy.

### v2

@@ A.1 Exchangeable Parameters and Values in CNNs and Embeddings
Normalization Layers. Normalization layers (NL) such as batch normalization [30] and layer normalization [4] can stabilize training. It applies a channel-wise scaling and bias. In CNNs, one normalization layer is usually inserted right before or after each activation function. We present an analysis for the latter case but the analysis also applies to the former case. Denote the NL functions as $\mathrm{NL}$ with parameters $\mathrm{nl}$, and $\mathrm{NL}_{i}$ denotes the NL applied to the $i$-th channel, parameterized by $\mathrm{nl_{i}}$. We instantiate $n=C_{2}$, $\zeta{}_{i}=W^{\prime}_{i\cdot}\oplus{}\mathrm{nl}_{i}\oplus{}W_{\cdot i}$, and $\xi{}_{i}=\mathrm{conv}(\mathrm{NL}_{i}(\mathrm{\sigma}{}(\mathrm{conv}(A,W^{\prime}_{i\cdot}))),W_{\cdot i})$.
@@
Embeddings Embedding dimensions are intuitively symmetric because embeddings are “distributed” representations as the relevant information is represented in many dimensions. We present the simple example Word2vec [41]. Let $m$ be the number of words, $n$ be the embedding dimension, $A$ be the embedding matrix for all the words, $M$ be the fully connected layer weight matrix (both of size $m\times n$), and $\mathrm{\sigma}{}$ be the softmax operation. The NN takes into input a word index $k$, and outputs $\mathrm{\sigma}{}(MA_{k}^{T})$. We instantiate $\zeta{}_{i}=A_{\cdot i}\oplus{}M_{\cdot i}$ and $\xi{}_{i}=A_{ki}M_{\cdot i}$. Alternatively, $\xi{}_{i}=A_{ki}$, which indicates that learned embedding dimensions are exchangeable.
@@ A.2 Evaluation Details
Dataset split. The test set labels of PIQA [6] are not published, so we use the validation set as the test set, and $10\%$ of the training set as the validation set. This is reasonable because the pretrained OPT model was not trained on PIQA training set. For CIFAR100 [36], we use a fixed split of the training set, with 90% samples used for training, 10% used as validation set, and use the default test set. We use the default dataset splits for ogbg-molhiv dataset [56].
Model architectures. We use the default OPT architecture from HuggingFace [55]. The GCN architectures follow Kipf and Welling [32]. We use the adapted CNNs [25] for CIFAR100. Specifically, the first layer and the pooling layer before the prediction head are adapted in size for the image size and class number in CIFAR100.
Obtaining trained models. We train image CNNs and GCNs locally using the training set, and used the pretrained weights for OPT [58] from HuggingFace [55]. For local training of ResNet18-BN, we use AdamW optimizer, $\times 10^{-2}$ learning rate, $16$ batch size, $10^{-4}$ weight decay, 1cycle learning rate scheduler, and 100 epochs. When statically pruning VGG11-BN on CIFAR10, we use the same training scheme in every pruning iteration to finetune the model after 5% parameters in all convolution kernels are set to zero. For local training of GCN, we use AdamW optimizer, $10^{-3}$ learning rate, $32$ batch size, $10^{-4}$ weight decay, 1cycle learning rate scheduler, and 80 epochs.
Optuna hyperparameter tuning. Across all models, we set the range of $\alpha$ as $[0,0.5]$ in StatsTest for all layers. The range of $T$ in Threshold is [30,0][-30,0] for CNNs, [1,0][-1,0] for GCN, [0.005,0][-0.005,0] for OPT. For statically pruned models, we additionally add a parameter $r$ in range $[0.1,0.5]$ to tune for each layer, which controls when ExPrune is disabled. When the ratio of the total FLOPs that can be potentially pruned for a channel’s computation to the overhead of ExPrune is below $r$, ExPrune is disabled. For each model, we provide a set of initial points to warm up Optuna’s surrogate model. Please see our code for details.